If you double Lisa's age and then subtract 6, you get Paul's age. 2 years ago, the sum of their ages was 29. How old is each person?
Answer by mananth(16949)

If you double Lisa's age and then subtract 6, you get Paul's age. 2 years ago, the sum of their ages was 29. How old is each person?
Let's assume Lisa's age to be x
double her age = 2x
subtract 6
2x-6 =paul's age
two years ago
lisa was x-2 years old
And paul was (2x-6)- 2
the sum of their ages was 29 2 years ago
x-2 + (2x-6) -2 =29
x-2 +2x-6 -2 =29
3x= 29+2+6+2
3x= 39
x= 39/3
x=13
lisa's age = 13
and Paul's age = 26
Ananth professional online tutor